OpenStudy (anonymous):

Solve : 3/8 - (1/5 + 3/4)

OpenStudy (anonymous):

\[\frac{ 3 }{ 8 } - (\frac{ 1 }{ 5 } + \frac{ 3 }{ 4 })\]

OpenStudy (anonymous):

Do parenthesis first

OpenStudy (anonymous):

19/20

OpenStudy (anonymous):

Find the common denominator of 5 and 4

OpenStudy (anonymous):

its 20

OpenStudy (anonymous):

convert 1/5 to 3/4 to them/20

OpenStudy (anonymous):

do its \[\frac{ 4 }{ 20 } + \frac{ 15 }{ 20 }\]

OpenStudy (anonymous):

so yes its 19/20

OpenStudy (anonymous):

now we need to \[\frac{ 3 }{ 8 } - \frac{ 19 }{ 20 }\]

OpenStudy (anonymous):

so find the common denominator again of 8 and 20

OpenStudy (anonymous):

140

OpenStudy (anonymous):

no its 80

OpenStudy (anonymous):

oh ok i see my mistake

OpenStudy (anonymous):

is the answer 57/80

OpenStudy (anonymous):

nevermind its not

OpenStudy (anonymous):

answer is 23/40

OpenStudy (anonymous):

its not..

OpenStudy (anonymous):

I got - 23/40

OpenStudy (anonymous):

how did you 23/40?

OpenStudy (anonymous):

yes